# SemEval-2024 Task 8: Weighted Layer Averaging RoBERTa for Black-Box Machine-Generated Text Detection

### 24 Feb 2024 v1

@@ Abstract
This document contains the details of the authors’ submission to the proceedings of SemEval 2024’s Task 8: Multigenerator, Multidomain, and Multilingual Black-Box Machine-Generated Text Detection Subtask A (monolingual) and B. Detection of machine-generated text is becoming an increasingly important task, with the advent of large language models (LLMs). In this document, we lay out the techniques utilized for performing the same, along with the results obtained.
[…]
@@ 1 Introduction
Language modeling is a foundational task in NLP, and encompasses learning of all the features that make up language. Different levels of linguistic information are stored in language models’ (LM) hidden states. This may include syntax, morphological features, phrasing, and so on. Our aim is to leverage this encoded information to help us discern machine-generated text.
@@ 2.1 Weighted Layer Averaging
The standard fine-tuning setup uses the [CLS] Representation of the last layer of RoBERTa. It has been shown that different layers of BERT-like models capture different levels of linguistic information [Rogers et al., 2020]. We believe that using just the last layer representation may discard some of the syntactic and lexical information, which could be crucial for the task of detecting machine generated text. We use the weighted sum of all the token representations, where each layer is assigned a corresponding weight, trained along with the downstream task, similar to ElMo [Peters et al., 2018].
@@ 3 Data
Data for the task was provided by the organizers. It is an extension of the M4 Dataset [Wang et al., 2023]. The name stands for multi-generator, multi-domain, and multi-lingual corpus for machine-generated text detection. As the name suggests, the dataset has been created with text from different generators spanning multiple domains. The data for subtask A and B follow the same format, consisting of source (such as Wikipedia), model (such as Dolly), label (such as Human), and the text to be classified. The data for subtask C contains text with a combination of human- and machine-generated text, and a label indicating the word index at which the split occurs.

### 09 Apr 2024 v2

@@ Abstract
This document contains the details of the authors’ submission to the proceedings of SemEval 2024’s Task 8: Multigenerator, Multidomain, and Multilingual Black-Box Machine-Generated Text Detection Subtask A (monolingual) and B. Detection of machine-generated text is becoming an increasingly important task, with the advent of large language models (LLMs). In this paper, we lay out how using weighted averages of RoBERTa layers lets us capture information about text that is relevant to machine-generated text detection.
$\dagger$\$$\dagger$\$footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Language modeling is a foundational task in NLP, and encompasses learning of all the features that make up language. Different levels of linguistic information are stored in language models’ (LM) hidden states. This may include syntax, morphological features, phrasing, and so on. [Rogers et al., 2021] Our aim is to leverage this encoded information to help us discern machine-generated text.
@@ 2.1 Weighted Layer Averaging
The standard fine-tuning setup uses the [CLS] Representation of the last layer of RoBERTa. It has been shown that different layers of BERT-like models capture different levels of linguistic information, the lower layers capture lexical information and word order,the middle layers capture syntactic information, and the higher layers capture semantic and task specific information [Rogers et al., 2020]. We believe that using just the last layer representation may discard some of the syntactic and lexical information, which could be crucial for the task of detecting machine generated text. We use the weighted sum of all the token representations, where each layer is assigned a corresponding weight, trained along with the downstream task, similar to ElMo [Peters et al., 2018]. Let $x_{0},x_{1},...x_{n}$ be the input sequence. Roberta generates the following hidden states.
$\texttt{RoBERTa}([x_{0},x_{1},\ldots,x_{n}])=H$
Where $H$ is a matrix consisting of hidden state vectors $\textbf{h}_{i}^{j}$ corresponding to the $j^{\text{th}}$ layer, and the $i^{\text{th}}$ token. $i=0$ represents the embedding layer output.
$H=\begin{bmatrix}\textbf{h}^{0}_{0}&\textbf{h}^{0}_{1}&\ldots&\textbf{h}^{0}_{n}\\ \textbf{h}^{1}_{0}&\textbf{h}^{1}_{1}&\ldots&\textbf{h}^{1}_{n}\\ \vdots&\vdots&\ddots&\vdots\\ \textbf{h}^{12}_{0}&\textbf{h}^{12}_{1}&\ldots&\textbf{h}^{12}_{n}\\ \end{bmatrix}$
The standard fine-tuning setup uses $\textbf{h}_{0}^{12}$ which corresponds to the [CLS] token and passes it through another Feed Forward Network to get the output class probabilities. We propose averaging all of the layer hidden states. The input y to the Feed Forward Network that produces the class probabilities is computed as follows.
$\textbf{y}=\frac{1}{12}\cdot\sum_{j=0}^{12}\frac{\lambda_{j}\sum_{i=0}^{n}\textbf{h}_{i}^{j}}{n}$
$\lambda_{j}$ is the layer weight assigned to the layer j. [$\lambda_{0},\lambda_{1},...\lambda_{12}$] are trained along with the classification task.
@@ 3 Data
Data for the task was provided by the organizers[Wang et al., 2024a]. It is an extension of the M4 Dataset [Wang et al., 2024b]. The name stands for multi-generator, multi-domain, and multi-lingual corpus for machine-generated text detection. As the name suggests, the dataset has been created with text from different generators spanning multiple domains. The data for subtask A and B follow the same format, consisting of source (such as Wikipedia), model (such as Dolly), label (such as Human), and the text to be classified. The data for subtask C contains text with a combination of human- and machine-generated text, and a label indicating the word index at which the split occurs.
@@ Appendix A Hyperparameters
$\dagger$\$$\dagger$\$footnotetext: The code used can be found in this repository: https://github.com/advin4603/
AI-Detection-With-WLA